A circular coil 15.0 cm in diameter and containing nine loops lies flat on the ground. The Earth's magnetic field at this location has magnitude 5.00×10-5T and points into the Earth at an angle of 53.0* below a line pointing due north. A 7.80-A clockwise current passes through the coil. Determine the torque on the coil.
Answer #1

given that

diameter d = 0.15 m

radius r = d/2 = 0.075 m

magnetic field B = 5*10^-5 T

current i = 7.8 A

from the relation

torque T =

where magnetic moment = NiA

T = (9*7.8*3.14*0.075^2 j)*5*10^-5*(cos53i-sin53j)

T = 6.19*10^-5*(-0.602 k)

T = torque = 3.726*10^-5 N-m (into the page)